An object 6 cm tall is placed 9 cm in front of a concave mirror that has a focal length of 18cm. Round your answers to 2 decimal places.  



1
Expert's answer
2022-01-13T09:21:34-0500

(a) We can find the image distance from the mirror equation:


"\dfrac{1}{d_o}+\dfrac{1}{d_i}=\dfrac{1}{f},""d_i=\dfrac{1}{\dfrac{1}{f}-\dfrac{1}{d_o}},""d_i=\dfrac{1}{\dfrac{1}{18\ cm}-\dfrac{1}{9\ cm}}=-18\ cm."

The sign minus means that the image is virtual and located behind the mirror.

(b) We can find the image height from the magnification equation:


"\dfrac{h_i}{h_o}=-\dfrac{d_i}{d_o},""h_i=-\dfrac{h_od_i}{d_o}=-\dfrac{6\ cm\times(-18\ cm)}{9\ cm}=12\ cm."

The sign plus means that the image is upright.